# Ev2R: Evaluating Evidence Retrieval in Automated Fact-Checking

### v1

@@ 3.2 Proxy-reference
Ev2R proxy-reference scorers use the veracity label $y$ as a proxy to evaluate the quality of the retrieved evidence $\hat{E}$ (see bottom right box in Figure 1). This approach is particularly relevant to rate evidence when no reference evidence is available but the veracity labels of the claims. For example, if a claims is labelled as false, we can use proxy scorers to evaluate if the retrieved evidence supports the claim or not.
Proxy fine-tuned scorer
The proxy trained scorer uses the DeBERTa language model (He et al., 2021) as backbone to predict the label $y$ serving as a proxy for evidence evaluation. We use a DeBERTa-large model that was previously trained on multiple benchmarks for the related task of natural language inference, specifically MNLI (Williams et al., 2018), Fever-NLI (Nie et al., 2019), Adversarial-NLI (Nie et al., 2020), LingNLI (Parrish et al., 2021) and WANLI (Liu et al., 2022). To fine-tune this model for the task of claim classification we use the datasets FEVER (Thorne et al., 2018a), VitaminC (Schuster et al., 2021), Hover (Jiang et al., 2020), and AVeriTeC (Schlichtkrull et al., 2023). Given a claim $c$, its veracity label $y$, and a retrieved evidence set $\hat{E}$ as input, the scorer predicts a score $s$ representing its confidence in the gold veracity label:
$s=\text{Scorer}(c,y,\hat{E})$
This score $s$ is derived by transforming the model-predicted output logits into probabilities using softmax. The resulting probability distribution allows the scorer to quantify its certainty in the different labels potentially available (e.g., support, refute, and not enough information). The scorer’s confidence in the gold label (e.g., $y=\text{support}$) is used as the output score $s\in[0,1]$. Hence, evidence contributing to high confidence in the gold label is rated higher accordingly.
LLM-based proxy scorer
@@
The LLM-based proxy scorer follows a similar approach. Instead of measuring the scorer’s confidence in the gold label $y$ based on the predicted probability distribution as described before, the scorer’s LLM is prompted with an input that consists of the instruction to predict a veracity label for the given retrieved evidence $\hat{E}$ and claim $c$. The predicted score $s$ corresponds to the following conditional probability:
$\text{Scorer}(c,y,\hat{E})=\log p(y\mid P(I,c,\hat{E}),\theta)$
The input prompt consists of a prompt template $P(.)$ that contains an instruction $I$ with three examples as demonstrations, as well as the input claim $c$ and evidence $\hat{E}$. The log-probability score $\log p(y\mid P(I,c,\hat{E}),\theta)$ corresponds to the models’s confidence in the label $y$ based on reasoning steps, as specified in the prompt template. We use the chain-of-thought prompting strategy (Wei et al., 2022), instructing the model to first output its chain of reasoning before predicting the final veracity label.
3.3 Reference-less
@@
The Ev2R reference-less scorers evaluate the retrieved evidence based only on the input claim (see top row in Figure 1). The scorer decomposes the claim $c$ into atomic facts $A_{c}$ and evaluates whether each fact $a_{c}\in A_{c}$ is addressed in the retrieved evidence $\hat{E}$ and either supported or refuted information is provided. This results in following score:
@@
Iterating over each atomic fact $a_{c}\in A_{c}$, the indicator function $\mathbb{I}$ returns $1$ if the atomic fact $a_{c}$ is either supported or refuted by the retrieved evidence $\hat{E}$ and 00 otherwise. This approach, similar to the reference-based scorer, is inspired by FactScore (Min et al., 2023). However, unlike the reference-based evaluation, which extracts evidence for both gold and retrieved evidence, this method decomposes the claim to evaluate the retrieved evidence without relying on any additional reference. The key advantage of this method is that it does not require any annotated data. Moreover, we compare our proposed evaluation framework to five reference-based baseline metrics that have been used in previous research for evidence evaluation: RougeL, BLEU, BLEURT, METEOR, and Hungarian Meteor.
4 Evaluation of Ev2R
This section provides an overview of the systematic evaluation approaches we conducted as part of the meta-evaluation of Ev2R scorers. Meta-evaluation consists of evaluation with the help of human evaluators (see Section 4.1) and automated evaluation using so-called checklist tests (see Section 4.2).
4.1 Evaluation with Human Ratings
We conduct a meta evaluation with human raters to assess how well the scorers’ predictions correlate with human judgements. We evaluate this based on five dimensions (detailed below) inspired by previous work on NLG metrics. [^3] See appendix for annotation instructions and screenshots of the evaluation platforms.
We asked annotators to rate retrieved evidence based on the following criteria while considering the provided reference evidence:
$(1)$ Coverage assesses how much of the reference evidence is covered by the retrieved evidence, whether the content, meaning, entities, etc., of the reference evidence are fully represented in the retrieved evidence.
$(2)$ Coherence captures whether the retrieved evidence is coherent, i.e., all sentences are connected sensibly and the evidence makes sense as a whole.
$(3)$ Repetition evaluates whether the retrieved evidence exhibits repetition of its content.
$(4)$ Consistency assesses if the retrieved evidence is logically consistent in itself.
$(5)$ Relevance measures how relevant the retrieved evidence is for the claim.
$(6)$ Verdict Agreement measures if the retrieved evidence results in the same verdict label as the reference evidence.
To assess the scorers, we considered two different data sources. First, we collected ratings for 100 randomly selected test data instances from the AVeriTeC test set (Schlichtkrull et al., 2023), along with the corresponding retrieved evidence from the AVeriTeC baseline system. Each sample was evaluated by two annotators, resulting in 200 annotations. The overall rating for each annotated sample and dimension was calculated as the average of both annotations. If the annotators disagreed on the verdict assigned to the retrieved evidence, a third annotator was assigned to the data instance, and majority voting was used to determine the final verdict. All human annotators were familiar with AFC research and were either computer science graduate students or postdocs.
The second source for evaluation was the AVeriTeC shared task submissions (Schlichtkrull et al., 2024). In collaboration with the organisers we obtained a random sample of evidence across all submitted systems, representing both high- and low-scoring systems. The participants were then asked to evaluate these samples. Thirteen participating teams manually evaluated thirty evidence samples each, five of which were gold-labeled samples annotated by the ourselves. These gold-labeled samples were included to aid in the post-processing of the collected annotations and to assess their quality. Excluding the gold-labeled samples, resulted in a final set consisting of $278$ evidence annotations.
Scorer Coverage Coherence Repetition Consistency Verdict Agree Avg
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
[…]
Ref-less .226 .254 .032 .138 .044 .077 .072 .067 .031 .035 .093 .134
Ref-based (prec) .458 .41 .304 .297 .148 .158 .147 .146 .262 .227 .248 .227
Ref-based (recall) .509 .530 .354 .372 .235 .26 .156 .167 .269 .262 .296 .306
Proxy-ref .114 .113 .122 0.131 .076 .094 .137 0.14 .499 .499 .199 .204
Prompt Scorer Gemini-Pro
Ref-less .459 .460 .262 .324 .138 .173 .353 .365 .333 .334 .309 .331
Ref-based (prec) .492 .435 .323 .329 .115 .119 .083 .089 .132 .111 .228 .217
Ref-based (recall) .578 .565 .309 .283 .210 .183 .015 .025 .257 .257 .273 .263
Prompt Scorer Gemini-Flash
Ref-less .347 .353 .183 .255 .124 .144 .283 .294 .348 .346 .257 .278
Ref-based (prec) .532 .473 .305 .310 .164 .168 .053 .054 .161 .137 .244 .228
Ref-based (recall) .619 .581 .325 .296 .168 .156 .147 .135 .226 .221 .297 .278
Prompt Scorer Llama 3.1
Ref-less .445 .425 .220 .319 .109 .156 .210 .225 .272 .250 .251 .275
Ref-based (prec) .178 .157 .048 .088 -.141 -.093 .091 .112 .233 .246 .082 .102
Ref-based (recall) -.028 .002 .036 .012 -.240 -.230 .125 .105 .148 .141 .008 .006
Trained Scorer
Ref-based .131 -.088 -.026 -.225 .182 .064 -.143 -.125 .207 .174 .076 -.020
Proxy-ref .271 .225 .164 .039 .057 -.013 .320 .262 .471 .485 .29 .253
Baselines
RougeL .35 .259 .07 -.061 -.005 -.077 .079 -.017 .056 .052 .091 .016
BLEU .281 .203 .253 .146 -.01 -.016 .063 .061 .144 .081 .133 .081
Meteor .369 .338 .084 .11 .17 .182 -.06 -.048 .08 .054 .114 .106
H-METEOR .136 .102 .093 .101 .043 -0.032 .113 .106 -.034 -.065 .039 .011
Table 1: Correlation between human-rated subset of AVeriTeC test and proposed scorers. We calculate correlation using the Spearman ($\rho$) and Pearson ($r$) correlation coefficients. Across all assessed dimensions (coverage, coherence, etc.), the proposed Ev2R scorers outperform the baselines, with prompt-based scorers yielding the highest average correlation scores.
4.2 Checklist Evaluation
Additionally, we evaluate the Ev2R scorers by assessing their sensitivity to automated perturbations in evidence data. Inspired by Ribeiro et al. (2020), we generated large-scale sets of so-called checklist tests by perturbing the claim and/or evidence. We included both semantics-preserving and semantics-modifying tests. For semantics-preserving tests, we change the claim and/or evidence such that the resulting evaluation score should not be affected, such as by introducing small typos. These tests assess the scorers’ robustness to semantically equivalent changes in the evidence data. On the other hand, we also create semantics-modifying tests by changing the evidence such that its meaning alters and subsequently scores should drop, for example, by removing a significant part of he evidence. To create tests with automated perturbations, we use the AVeriTeC test set as the basis.
Our tests focus on the seven different dimensions (see Table 4 in the appendix), the first two categories alter the semantics of the initial evidence, while the remaining ones preserve the evidence’s meaning.
4.3 Backbone models
In Section 3, we introduce a framework for evidence evaluation. As highlighted earlier, various language models can be integrated as backbone models within this framework. In this work, we assess the framework using five different LLMs.
@@ 5 Results and Discussion
In this section, we discuss the results obtained by evaluating the evidence scoring framework using three different datasets. First, we evaluate all Ev2R scorers based on the AVeriTeC subset, which was scored by human annotators (see Table 1). Second, we use human-evaluated system predictions from the AVeriTeC shared task (see Table 2). Finally, we compare all Ev2R scorers and baselines against each other based on the automatically generated adversarial tests (see Table 3). Following Fu et al. (2023), we measure the correlation between the Ev2R scorers and human ratings using Spearman ($\rho$(Spearman, 1987) and Pearson correlation coefficients ($r$(Pearson, 1896).
Ev2R scorers versus traditional metrics
@@
Comparison: Reference-based, reference-less, and proxy-reference
Among all prompt scorers (reference-based, reference-less, and proxy-reference), the reference-based Ev2R scorers consistently achieve higher correlation scores than the reference-less ones across the evaluated datasets. This is expected, as the availability of gold-annotated evidence, such as in the AFC datasets, aids in understanding and assessing predicted evidence. However, when compared to baseline models that rely on reference evidence for scoring, the Gemini Pro-based reference-less scorer also shows promising results.

### v2

@@ 3.2 Proxy-reference
The proxy-reference scoring component of Ev2R uses the verdict label $y$ as a proxy to assess the retrieved evidence $\hat{E}$. As shown in the right box of Figure 1, this component compares the predicted verdict label to the reference label, i.e., refuted. Proxy-reference scoring uses a fine-tuned DeBERTa language model (He et al., 2021) as backbone to predict the verdict label $y$ serving as a proxy for evidence evaluation.
The proxy-reference scoring component builds on a DeBERTa-v3 model (He et al., 2021) finetuned to predict verdict labels for claim-evidence pairs. We initialised the scorer from a checkpoint pretrained on a diverse set of NLI datasets, i.e., MNLI (Williams et al., 2018), FEVER-NLI (Nie et al., 2019), Adversarial-NLI (Nie et al., 2020), LingNLI (Parrish et al., 2021), and WANLI (Liu et al., 2022), which provide a foundation for assessing factual inference. To adapt the model for evidence evaluation in the context of automated fact-checking, we fine-tuned it further on AFC datasets, including the training sets of FEVER, VitaminC, HOVER, and AVeriTeC. The training was conducted using the HuggingFace Trainer API, mixed-precision (fp16) and an initial learning rate of $1\text{e}^{-6}$. The model was trained for three epochs, a warmup phase of fifty steps, and evaluated every $10k$ steps using the micro-F1 score on held-out evidence sets to select the best checkpoint.
Given a claim $c$, a retrieved evidence set $\hat{E}$, and a reference label $y\in\mathcal{Y}$, the proxy scoring component $\mathcal{S}_{\text{proxy}}$ assigns a confidence score $s_{\text{proxy}}$ to the label $y$, reflecting the model’s confidence in the predicted verdict:
$s_{\text{proxy}}$ $=\mathcal{S}_{\text{proxy}}(c,y,\hat{E})$
$=\text{softmax}(\mathbf{z})_{y}=\frac{e^{z_{y}}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{z_{y^{\prime}}}}$
where $\mathbf{z}$ is the vector of logits over all possible veracity labels $\mathcal{Y}$. The score $s_{\text{proxy}}\in[0,1]$ is computed as the softmax of the logits $\mathbf{z}$, where the index $y$ corresponds to the reference label, and the softmax function transforms the logits into a probability distribution over the possible labels. The resulting softmax distribution quantifies the model’s confidence.
3.3 Final Weighted Scorer
To compute the final Ev2R score, we use the outputs of both previously introduced components: the reference-based scorer $\mathcal{S}{\text{ref}}(\hat{E},E)$, which yields precision and recall scores ($s{\text{prec}}$ and $s_{\text{recall}}$), and the proxy scorer $\mathcal{S}_{\text{proxy}}(c,y,\hat{E})$, which estimates the confidence in the label $y$. We first compute the $F_{1}$ score from the precision and recall outputs of $\mathcal{S}_{\text{ref}}$. Then we calculate the final weighted score as a combination of the resulting $F_{1}$ score and the $s_{\text{proxy}}$. Specifically, we use a weighting factor $\alpha=0.5$ to equally balance both components:
$s_{\text{{\color[rgb]{0,0.3984375,0.80078125}Ev\textsuperscript{2}}}\text{{\color[rgb]{0.80078125,0.3984375,0}R}}}=\alpha\cdot\left(\frac{2\cdot s_{\text{prec}}\cdot s_{\text{recall}}}{s_{\text{prec}}+s_{\text{recall}}}\right)+(1-\alpha)\cdot s_{\text{proxy}}$
This ensures that the final score reflects both the factual correctness and completeness of the retrieved evidence with respect to the reference, as well as its alignment with the verdict.
4 Evaluation of Ev2R
This section provides an overview of the systematic evaluation approaches conducted as part of the meta-evaluation of Ev2R, which involves three different data sources. First, using adversarial tests generated through perturbing evidence data (see Section 4.1). Second, evidence obtained from AFC systems submitted to the AVeriTeC shared task competition (Schlichtkrull et al., 2024b) in early 2024. (see Section 4.2). Third, evidence pairs extracted from the AFC benchmarks FEVER (Thorne et al., 2018a) and VitaminC (Schuster et al., 2021). We discuss the data sources in more detail below. Finally, in Section 4.4, we give an overview of baseline scorers and models used for evaluation.
4.1 Evaluation with Adversarial Tests
We evaluate the scorers by assessing their robustness against adversarial perturbations, including both semantics-altering and semantics-preserving changes. Specifically, we perform adversarial stress tests, including both semantics-altering and semantics-preserving perturbations, such as redundancy, noise insertion, and completeness alterations. Inspired by Ribeiro et al. (2020), we generated large-scale test sets by perturbing either the claim or evidence data. We included both semantics-preserving and semantics-modifying tests (see Table 6 in the appendix). For semantics-preserving tests, we change the claim and/or evidence such that the resulting evaluation score should not be affected, for example, by introducing small typos. These tests assess the scorers’ robustness to semantically equivalent changes in the evidence data. On the other hand, we also create semantics-modifying tests by changing the evidence such that its meaning alters and subsequently scores should drop, for example, by removing a significant part of the evidence. To create tests with automated perturbations, we use the AVeriTeC test set as the basis.
4.2 Evaluation with Human Ratings
Data source
The AVeriTeC dataset involves retrieving evidence from the web and assessing the veracity of real-world claims previously checked by professional fact-checkers. It addresses limitations in prior AFC benchmarks through focusing on real-world claims and annotated evidence in question-answer format. The AVeriTeC shared task attracted a total of $21$ submissions of fact-checking systems (Schlichtkrull et al., 2024a). Participating systems were required to provide URLs for their retrieved evidence as well as verdict labels categorized as supported, refuted, not enough evidence, or conflicting evidence/cherrypicking. The organisers provided a public knowledge store with documents to avoid reliance on costly search APIs.
In collaboration with the AVeriTeC organisers, we obtained a subset of evidence across all submitted systems, equally representing both high- and low-scoring systems. After ranking the systems according to their performance on the shared task, we categorise them in four performance groups each containing approximately $25\%$ of submitted systems to allow comparative analysis: Low score (below $0.16$), mid-low score (between $0.16$ and $0.22$), mid-high score (between $0.22$ and $0.435$), and high score ($0.435$ or higher). This classification enabled systematic examination of system performance across different scoring levels. Using stratified sampling, we selected $560$ claims to ensure an even distribution across the dataset.
Human ratings and agreement
The shared-task participants were then asked to evaluate these samples. [^3] Find details on the annotation platform in the appendix. Thirteen participating teams manually annotated thirty evidence samples each, five of which were gold-labeled samples annotated by ourselves. For the final set, we only considered annotation submissions of a team, if the gold-labeled samples were correctly annotated, resulting in a final set consisting of $278$ evidence annotations. To measure the agreement between annotators, we used two methods. First, we calculated the agreement between the categorical verdict labels that annotators selected as the first step in the annotation process. Therefore, we used all $278$ annotated samples after filtering as specified above, and obtained a Krippendorff’s Alpha (K-$\alpha$(Krippendorff, 2004) $0.732$ and a Fleiss’ Kappa (Fleiss, 1971) (F-$\kappa$) of $0.727$ both indicating substantial agreement. Second, for the rating dimensions which are annotated with numeric values between one and five, we calculated the standard deviation (std) among the rating, resulting in std between $0.804$ and $1.318$ (see Table 1).
Rating Dimension IAA method Score
Verdict agreement Fleiss-$\kappa$ 0.727
Verdict agreement Krippendorff-$\alpha$ 0.732
Coverage Std 1.119
Relevance Std 0.804
Table 1: Inter-annotator agreement (IAA) scores for human ratings using Fleiss-$\kappa$/Krippendorff-$\alpha$ for categorical (i.e., verdict agreement) and standard deviation (Std) for continuous ratings.
Evaluation dimensions
With the help of human raters, we assessed how well the scorers correlate with human judgments. For this evaluation, we considered three dimensions:
$(1)$ Coverage assesses how much of the reference evidence is covered by the retrieved evidence, whether the content, meaning, entities, etc., of the reference evidence are fully represented in the retrieved evidence, ensuring that the retrieved evidence covers all aspects of the reference evidence necessary for the evaluation. Partial evidence may overlook certain details and result in incomplete or incorrect conclusions.
$(2)$ Relevance measures the relevance of the evidence retrieved for the claim, ensuring that the evaluated evidence addresses the claim and that minimal unrelated information is included.
$(3)$ Verdict Agreement measures if the retrieved evidence results in the same verdict label as the reference evidence. Discrepancies in verdicts can indicate issues with the retrieved evidence such that a follow-up review can be performed to ensure accuracy in the fact-checking process.
While coverage can penalise potentially valid evidence, this only occurs when such evidence is included at the expense of the gold/reference evidence. As long as the gold evidence is fully covered, the coverage score remains high, even if additional content is included. This differs from the relevance score we introduced above, which is more flexible: it does not penalise the inclusion of content different to the gold evidence, as long as the content is relevant for verifying the claim. We report both alignment with relevance and coverage, with the latter being relevant for the recall-oriented automatic Ev2R evaluation.
4.3 Evaluation with additional benchmarks
In addition to the AVeriTeC shared task data, we use the large-scale AFC benchmarks FEVER (Thorne et al., 2018a) and VitaminC (Schuster et al., 2021). FEVER is a benchmark designed for automated claim verification against textual sources from Wikipedia. It contains $185.4k$ claims generated through modifying sentences from Wikipedia and classified as supported/refuted/not enough information. VitaminC is an AFC benchmark designed for robust claim verification. It consists of approximately $400k$ pairs of claims and evidence pairs derived from Wikipedia revisions. The aim was to create contrastive pairs of nearly identical evidence, differing only by subtle factual details, such that one evidence set supports the claim and the other refutes it, or vice versa. For scorer assessment, we created tuples consisting of a claim $c_{i}$, a reference evidence set $E_{ref}$, predicted evidence set $E_{pred}$ from the multiple evidence sets available for individual claims within the FEVER and VitaminC test datasets. If both evidence sets resulted in the same verdict label, we assigned a verdict agreement score of $1$; otherwise, we assigned a score of 00. The resulting data were used for evaluations presented in Table 2.
4.4 Baselines
To validate the effectiveness of Ev2R, we compare it to three types of baseline scorers: $(i)$ the reference-based BLEURT scorer, which compares retrieved evidence to reference evidence; $(ii)$ proxy-reference scorers, which assesses evidence through predicted verdict labels; and $(iii)$ reference-less scorers, which rely only on the claim and retrieved evidence. [^4] Prompts and further details are given in the appendix.
$(i)$ Reference-based baseline scorers
As a reference-based baseline, we adapt a scorer built on BLEURT (Sellam et al., 2020), originally developed for NLG tasks such as summarization. BLEURT uses BERT as its backbone and is trained on millions of synthetically generated sentence pairs to capture semantic similarities between texts. To evaluate the retrieved evidence against reference evidence, the BLEURT-based scorer compares both evidence sets and gives a similarity score. The final aggregated score $s$ is calculated by averaging instance-level scores across all evaluated samples.
Given a set of retrieved evidence $\hat{E}=\{\hat{e}_{1},\hat{e}_{2},\ldots,\hat{e}_{n}\}$ and reference evidence $E=\{e_{1},e_{2},\ldots,e_{n}\}$, the scorer computes:
$s=\text{BLEURT}(E,\hat{E})$
where $s\in[0,1]$, with higher values indicating greater semantic similarity between the evidence.
We fine-tune BLEURT for evidence evaluation using the Pytorch version of the original BLEURT model [^5] Elron/bleurt-base-512 and the HuggingFace Trainer API over 5 epochs, with a batch size of 4 and learning rate set to $1\text{e}^{-5}$. The model is trained on claim-evidence input pairs from AFC datasets including FEVER (Thorne et al., 2018a), VitaminC (Schuster et al., 2021), and HOVER (Jiang et al., 2020). These datasets provide multiple annotated evidence sets ${E_{1},E_{2},\ldots,E_{n}}$ per claim, enabling the construction of diverse similarity pairs. we generate input-label pairs where positive examples (similarity score $s=1$) are created using two distinct evidence sets aligned to the same claim or based on matching claim-verdict pairs (e.g., both supporting the claim). Negative examples ($s=0$) are created by pairing evidence from different claims or by contrasting verdict-aligned pairs (e.g., one supporting, one refuting).
Additionally, we include metrics such as RougeL, BLEU, METEOR, and Hungarian METEOR as reference-based baselines since they have been previously also used in AFC for evidence evaluation (Schlichtkrull et al., 2024b). They were originally developed for NLG tasks and compare retrieved ($\hat{E}$) with reference evidence ($E$) based on surface level or semantic similarity.
$(ii)$ Proxy-reference LLM baseline
@@ 4.4 Baselines
As a proxy-reference baseline, we propose an LLM-based scorer that estimates the likelihood of the correct veracity label $y$ given the claim $c$ and the retrieved evidence $\hat{E}$. Instead of relying on a predicted probability distribution as in trained classifiers, we prompt the LLM directly with an instruction to predict the veracity label and use its log-probability for the correct label as the proxy score. The scorer computes a score $s$ as follows:
$s=\log p(y\mid P(I,c,\hat{E}),\theta)$
where $P(I,c,\hat{E})$ is the prompt including an instruction $I$, the input claim $c$, and the retrieved evidence $\hat{E}$, and $\theta$ denotes the parameters of the LLM. We apply chain-of-thought prompting (Wei et al., 2022) and instruct the model to generate intermediate reasoning steps before selecting a final label. The prompt includes few-shot examples and reasoning steps as demonstrations to improve prediction quality. For the proxy-reference scorer, we use DeBERTa as the backbone model. Our choice is motivated by its strong performance on claim verification and natural language inference (NLI) tasks under fine-tuned settings, as well as its compatibility with our available compute resources (one A100100100100 GPU).
$(iii)$ Reference-less baseline
@@
As a reference-less baseline, we instruct an LLM-based scorer evaluate the retrieved evidence $\hat{E}$ based solely on the input claim $c$, without relying on any reference evidence. The scorer decomposes the claim into atomic facts $A_{c}$ and checks whether each fact $a_{c}\in A_{c}$ is addressed, for example, supported or refuted, by the retrieved evidence. The score is computed as follows:
@@
This approach, similar to the Ev2R reference-based component, is inspired by FactScore (Min et al., 2023). However, unlike the reference-based evaluation, which extracts evidence for both gold and retrieved evidence, this method decomposes the claim to evaluate the retrieved evidence without relying on any additional reference. [^6] The reference-less evaluation prompt is given in Figure 2 in the appendix.
Backbone models
@@ 5 Results and Discussion
In this section, we discuss the results obtained with Ev2R using three different data sources. First, human-rated system predictions from the AVeriTeC shared task (see Table 4). Second, the FEVER and VitaminC test sets for which we obtained pairs of evidence following the approach outlined in Section 4.4. Third, adversarial tests as described in  Table 5. Following Fu et al. (2023), we measure the correlation between the scorers and human ratings using Spearman ($\rho$(Spearman, 1987) and Pearson correlation coefficients ($r$(Pearson, 1896).
Scorer VitaminC FEVER AVeriTeC Avg
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $|\rho|$ $|r|$
Reference-based Baselines
BLEURT -.065 -.072 -.003 -.000 .151 .098 .073 .057
RougeL .030 .022 -.010 -.009 .070 .097 .037 .043
BLEU .030 .026 -.006 -.003 .166 .137 .067 .055
Meteor .030 .021 -.010 -.003 .152 .159 .064 .061
H-METEOR .001 -.001 -.027 -.016 -.029 -.019 .019 .012
Reference-less Baselines
@@
Gemini-Pro .034 .033 .415 .412 .262 .273 .237 .239
Gemini-Flash .038 .045 .434 .426 .230 .247 .234 .239
Llama 3.1 .024 .034 .293 .308 .215 .208 .177 .183
Ev2R Weighted Scorer
GPT-4o / Proxy-ref .276 .318 .366 .360 .500 .488 .381 .389
Gemini-Pro / Proxy-ref .295 .334 .349 .346 .472 .464 .372 .381
Gemini-Flash / Proxy-ref .336 .346 .170 .240 .129 .197 .292 .294
Llama / Proxy-ref .272 .322 .026 .227 .324 .307 .243 .273
Table 2: Correlation on the Verdict Agree category calculated using the Spearman ($\rho$) and Pearson ($r$) correlation coefficients for the datasets VitaminC, FEVER, and AVeriTeC (shared task subset). Averages are computed over the absolute values of each metric per row. Highest scores per column are colored blue and second-highest values highlighted with brown color.
Scorer VitaminC FEVER AVeriTeC Avg
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $|\rho|$ $|r|$
Reference-based (Precision)
[…]
Gemini-Pro .222 .216 .139 .118 .237 .177 .199 .170
Gemini-Flash .202 .198 .106 .079 -.290 -.205 .199 .161
Llama 3.1 -.037 -.036 .110 .109 .216 .210 .121 .118
Reference-based (Recall)
[…]
Gemini-Pro .253 .246 .176 .143 .285 .278 .238 .222
Gemini-Flash .232 .227 .032 .000 -.190 -.039 .151 .089
Llama 3.1 .020 .025 -.011 -.012 .128 .121 .053 .053
Reference-based ($F_{1}$)
[…]
Gemini-Pro .237 .230 .155 .129 .259 .216 .217 .193
Gemini-Flash .216 .212 .049 .000 -.230 -.066 .172 .115
Llama 3.1 .087 .164 -.024 -.027 .161 .154 .074 .073
Proxy-reference
Proxy-ref .456 .479 .290 .480 .487 .459 .411 .473
Table 3: More detailed scores for both Ev2R components: the reference-based part (using different LLMs as backbone models) and the DeBERTa-based proxy-reference component.
General insights across datasets
[…]
Ref-less Baselines Proxy-ref Ref-based Baselines Ev2R
GPT-4o Gem-Pro Gem-Flash Llama DeBERTa RougeL BLEU Meteor H-Met GPT-4o Gem-Pro Gem-Flash Llama
COV ($\rho$) .237 .287 .275 .297 .338 .150 .236 .229 .005 .321 .341 .221 .328
COV ($r$) .261 .296 .287 .286 .348 .169 .184 .240 -.024 .323 .326 .299 .336
REL ($\rho$) .292 .203 .287 .227 .298 .086 .107 .062 .008 .297 .278 .403 .224
REL ($r$) .360 .263 .333 .250 .374 .099 .079 .076 .003 .332 .315 .404 .277
Table 4: Correlation between human-rated AVeriTeC shared task submissions and proposed scorers for COVerage and RELevance dimensions. Highest scores per dimension are colored blue and second-highest values highlighted with brown color.
[…]
Ev2R weighted score versus traditional metrics
Across all datasets, i.e., VitaminC, FEVER, and AVeriTeC, the Ev2R scorer consistently outperforms traditional reference-based baselines we assess, including ROUGE-L, BLEU, METEOR, and H-METEOR. The performance gap is particularly prevalent on AVeriTeC, which includes real-world claims and evidence retrieved from the web. This setting is notably more complex than earlier benchmarks like FEVER and VitaminC. In such cases, traditional metrics that rely on surface-level or shallow semantic similarity often fail to generalize. These findings are supported by the insights we gain from the adversarial tests in Table 5, where we assess all baselines against both both Ev2R components, the proxy reference (last column) and reference-based (prec/recall) components. On all semantics-preserving tests, both Ev2R components yield robustness performance.
Ev2R: Balancing proxy and reference-based evidence evaluation
While the proxy-reference component of Ev2R helps stabilize scores on academic datasets with short claims and evidence (e.g., FEVER, VitaminC in Table 2), relying solely on proxy signals proves insufficient for more complex real-world scenarios such as AVeriTeC (see Table 2). Prior work (McCoy et al., 2019) has shown that NLI models often exploit shallow syntactic heuristics such as lexical overlap or subsequence matching that can lead to correct predictions for the wrong reasons. This motivated the need for a balanced evaluation approach that combines proxy-reference with reference-based assessment and offers a more reliable and interpretable measure of evidence data. Moreover, for relevant assessment dimensions such as Coverage of reference evidence and Relevance of retrieved evidence, the reference-based component of Ev2R yields higher correlation with human assessment.
Dataset Reference-less Baselines Ev2R Proxy-ref
METEOR ROUGE BLEU BLEURT GPT4o Prec Rec Proxy-comp. DeBERTa
Semantics-altering tests
Completeness -30.56 -53.19 -77.87 -143.70 -91.76 -21.2 -59.5 -39.35 -82.68
Random shuffle -32.32 -89.94 -94.20 -1.02 0.55 -6.8 -6.0 -32.22 -92.32
Average -31.44 -71.57 -86.04 -72.36 -45.61 -14.0 -32.75 -35.79 -87.50
Semantics-preserving tests
Invariance contraction -0.53 -0.97 -1.35 -1.65 0 -0.8 -0.7 -0.85 0.02
Invariance num2text -8.34 -10.24 -14.10 -22.13 -3.85 0.1 0 -4.02 -35.64
Invariance text2num -0.04 -0.90 -1.15 -0.81 -4.40 0 -0.3 -1.59 0.01
Invariance synonyms -8.22 -36.64 -51.83 -51.00 -2.75 -11.6 -11.3 -15.35 -85.71
Redundancy sent -26.16 -46.52 -56.15 -80.45 3.30 0 -0.1 -5.04 -67.95
Redundancy words -5.73 -10.68 -32.69 -31.90 -2.75 -1.0 0.2 -2.89 -52.55
Fluency -11.52 -23.33 -29.26 -41.69 -2.20 -4.3 -3.9 -2.6 -57.12
Noise -31.82 -31.81 -31.81 -21.20 -1.65 -28.7 -2.2 2.89 -62.42
Argument structure -0.01 0 0 0 1.65 0 0 -0.74 -28.93
Average -10.33 -18.43 -23.94 -27.87 -15.04 -5.16 -2.9 -3.78 -39.79
Table 5: Results obtained for adversarial tests as score difference (in %) between initial evidence and manipulated evidence. The evaluated Ev2R reference-based scoring uses $GPT4o$. Overall, the reference-based scoring and the proxy-reference scoring components demonstrate the most robust performance on adversarial tests, together with the trained reference-based baselines, as evidenced by small performance drops on semantics-preserving tests. Best scores per evaluation category are colored blue and second-best values highlighted with brown color.